find the general solution of the differential equation

55y''''+224y''+225y=0
Answer #1

Given differential equation

55y''''+224y''+225y=0

The characteristic equation is:

55r^4+224r^2+225=0

This can be written as:

55(r^2+\frac{112}{55})^2-\frac{169}{55}=0

The roots are:

(5r^2+9)(11r^2+25)=0

Therefore,

r^2=-\frac{9}{5},-\frac{25}{11}

r=\frac{3i}{\sqrt{5}},\frac{5i}{\sqrt{11}}

The general solution is:

y(t)=C_1\cos(\frac{3}{\sqrt{5}})t+C_2\sin(\frac{5}{\sqrt{11}})t